OpenStudy (anonymous):

What is the factored answer to 4x^2+8x-4

OpenStudy (anonymous):

Factor out the 4

OpenStudy (anonymous):

4(x^2+2x-1). Then factor the ( )

OpenStudy (anonymous):

I know you have to factor out the four but then I am confused how to factor the rest

OpenStudy (anonymous):

Isn't think u can factor any further.

OpenStudy (anonymous):

alright thank you

OpenStudy (ranga):

x^2+2x-1 = 0 does not have rational roots so you can leave it at 4(x^2 + 2x - 1)

OpenStudy (anonymous):

ok well what I am doing must b wrong so can someone help me solve this rational thing 

OpenStudy (anonymous):

??

OpenStudy (anonymous):

When u divide u must flip the fraction after the division sign. Factor everything. 3(x-4)(x+1). (X+3)(x+2) ------------------------------- 4(x^2-2x-1). (X-4)(x+3)

OpenStudy (anonymous):

= 3(x+1)(x+2) --------------------- 4(x^2-2x-1)

OpenStudy (anonymous):

I cross cancelled what I could to get the answer.

OpenStudy (ranga):

\[\Large \frac{ 3(x+1)(x+2) }{ 4(x^2+2x-1) }\]